# Alternative ranking measures to predict international football results

### v2

@@ 4 Applications
The selection of the training and test sets is crucial and is likely to influence the predictions. We address this by employing an iterative training approach. Our goal-based statistical models and result-based machine learning algorithms are trained on a continuously updated dataset encompassing international matches from $2018$ to $2023$. The matches vary from the FIFA World Cups through the UEFA Euro Championships to normal friendly matches. The data excludes Olympic Games and matches in which at least one of the teams was the national B-team or a U-23 lineup. This approach allows for the incorporation of recent results, potentially improving predictive performance for two distinct scenarios: the group stage and the knockout stage of this two tournaments.
We evaluate the predictive performance of three dynamic goal-based Poisson models, implemented using the footBayes package (Egidi and Palaskas.,, 2022), and three result-based machine learning techniques provided by the caret package (Kuhn,, 2022), both implemented in R, along the lines described in Section 3. Furthermore, we incorporate additional historical information from both the FIFA ranking and the Bayesian BTD derived ranking using the R bpcs package (Issa Mattos and Martins Silva Ramos,, 2020). Specifically, the integration as additional predictor of the Bayesian BTD derived ranking involves the following steps
  1. 1.
    Fit the Bayesian BTD model as described in (5).
  2. 2.
    Compute the posterior median for each team’s log-strength parameter.
  3. 3.
    Add the difference in the posterior median of the competing teams as additional predictor in both the goal-based models and result-based algorithms.
To ensure a more comparable analysis, both rankings - the FIFA and the BTD - are normalized using the scaled median absolute deviation (MAD) normalization
@@
where $\mathbb{M}(\cdot)$ is the median. Furthermore, to assess the predictive performance of the models described in Section 3, we employed the Brier score (Brier,, 1950) as recommended by Spiegelhalter and Ng, (2009). It is essentially a mean squared error for forecasts, with values between 00 and $2$, where a lower score indicates greater model predictive accuracy. A common formulation is
@@
The World Cup presents an interesting case study due to the diverse range of National teams with heterogeneous strengths. We specifically investigate the performance of teams in a structured environment, such as the group stage, and contrast it with the dynamic and high-pressure setting of the knockout stage, where single-elimination games can significantly impact teams behaviour. This subsection delves into how we employ both goal-based statistical models and machine learning algorithms to forecast match outcomes throughout the tournament, assessing their predictive performance adding the historical information from the ranking systems outlined previously. Notably, we consider the FIFA ranking published just before the World Cup took place, available at https://inside.fifa.com/fifa-world-ranking/men?dateId=id13869.
@@
The top panel of Figure 1 presents scatterplots comparing the ranks of competing teams in both the group and knockout stages under the two ranking systems. Points above the dashed line represent matches where the “away” team had a higher rank than the “home” team, while points below the line indicate the opposite - we note that the terms ‘home’ and ‘away’ do not mean anything relevant in a international competition, where there are just one or two hosting teams. As expected, both the ranking systems reveal greater variability in team ranks during the group stage compared to the knockout stage. This is because teams in the group stage are typically more heterogeneous in strength level. As the tournament advances to the knockout stage, the remaining teams become increasingly similar in ability, leading to less variation in rank. This pattern is even more evident in the bottom panel of Figure 1, which displays boxplots of rank differences between competing teams for each ranking system across the two stages. The boxplots further reveal that the Bayesian BTD ranking system exhibits more variability than the FIFA ranking during the group stage. In contrast, during the knockout stage, rank differences variability under both ranking systems are more concentrated, indicating a closer similarity in team strengths.
@@
Figure 1: 2022 FIFA World Cup. The top panel displays scatterplots comparing the ranks of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of rank differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two World Cup stages.
@@
The main appeal of these models lies in their ability to predict the outcome of football matches. Figure 2 illustrates the predictive accuracy of dynamic Poisson models and machine learning algorithms, measured by the Brier Score, across both group and knockout stages. The goal-based statistical models perform slightly better than the result-based methods during the group stage. In contrast, the machine learning approaches show better predictive accuracy in the knockout stage, where results may be less predictable. A similar trend is seen with ranking systems. While the FIFA ranking system shows a marginally better predictive accuracy in the group stage, the Bayesian BTD ranking demonstrates overall better performance during the knockout stage. This suggests that the Bayesian BTD ranking is particularly apt at predicting outcomes when teams have comparable strength, which is often the case in the latter stages of the competition.
@@
Figure 2: 2022 FIFA World Cup. Brier score for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two World Cup stages.
@@
In this section, we fit the considered statistical models and machine learning algorithms to the data from the most recent CAF Africa Cup of Nations (AFCON) tournament held in Ivory Coast. Notably, the AFCON competition differs from the FIFA World Cup in that the participating teams tend to be more similar in terms of overall strength even during the group stage, making it a compelling case for analyzing the effectiveness of the Bayesian BTD ranking system. In order to conduct this analysis, we use as training set the data from matches played throughout the period 2018 to the end of 2023. Furthermore, the Bayesian BTD model was executed within this same period to generate team rankings. In addition, the FIFA ranking employed correspond to those published on December 21st, available at https://inside.fifa.com/fifa-world-ranking/men?dateId=id14233.
@@
In this scenario, the Bayesian BTD derived rankings exhibit less variability compared to the FIFA ranking, in both the group and knockout stages, as illustrated in the top panel of Figure 3. Furthermore, the boxplots displayed in the bottom panel of Figure 3 show a significant reduction in variance for both ranking systems as we move from the group stage to the knockout stage. However, contrary to what observed in the World Cup, the Bayesian BTD derived ranking exhibit lower variance compared to the FIFA ranking in the group stage. This may indicate that the Bayesian BTD ranking more accurately reflect the inherent similarity in team strengths within this tournament stage.
@@
Figure 3: 2023 CAF Africa Cup of Nations. The top panel displays scatterplots comparing the ranks of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of rank differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two Africa Cup of Nations stages.
@@
As reported in Figure 4, all the machine learning algorithms presents similar prediction performance in the group stage of the AFCON. However, random forests show the weakest performance in the knockout stage. Furthermore, the other machine learning algorithms and statistical models exhibited similar levels of predictive accuracy in both stages, with statistical models performing slightly better in the group stage. In particular, the inclusion of this alternative ranking system generally improved the predictive performance of most of the statistical models evaluated in the knockout stage.
@@
Figure 4: 2023 CAF Africa Cup of Nations. Brier score for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two Africa Cup of Nations stages.
@@ 5 Discussion
This paper investigates the potential improvement in predictive performance of statistical goal-based methods and machine learning result-based algorithms when a ranking system is incorporated as an additional predictor variable. We analyze data from the recent 2022 FIFA World Cup in Qatar and the 2023 CAF Africa Cup of Nations in Ivory Coast. Specifically, we explore the effectiveness of a Bayesian Bradley-Terry-Davidson derived ranking system in enhancing prediction accuracy compared to the well-established FIFA ranking system. We compare the performance of these two ranking systems across different tournament stages to identify their relative strengths and weaknesses in predicting match outcomes.
@@
While both the FIFA ranking and the Bayesian BTD derived ranking provide valuable information for predicting outcomes, their effectiveness depends on the stage of the tournament. The FIFA ranking tends to be more accurate during the group stages of the 2022 World Cup, which features a more heterogeneous team strengths. Conversely, the Bayesian BTD derived ranking is particularly effective in the group stage of the 2023 AFCON, and it also presents slightly better predictive performances in the knockout stages of both the World Cup and the AFCON. These stages typically exhibit smaller differences in team strengths. Consequently, this result suggests that the Bayesian BTD model effectively captures shifts in team strengths, making it especially valuable in tournaments where competing team are similar, such as the AFCON, or in stages where the teams exhibit comparable strengths.
@@
It is important to note that while the Bayesian BTD derived ranking is a valuable alternative, it is more computationally intensive compared to the standard FIFA ranking. Specifically, this approach involves a two-step procedure. First, the Bayesian BTD model needs to be computed to derive the ranking, and only then the derived ranking can be used as additional predictor in the models being considered.
@@
However, the potential to enhance the accuracy and predictive performances of these models remains significant. Further development could involve refining the Bayesian BTD model to include additional variables that impact match outcomes. These could include the overall market value of the players involved in a specific team, the number of Champions League players, an indicator of the hosting country or of the teams in its neighbourhood. Even economic variables such as the GDP per capita or the national population size may be interesting. Furthermore, the implementation of a dynamic methodology, that enables the continuous adjustment for fluctuations in team strength over the course of a season or tournament, could potentially result in enhanced prediction accuracy.
@@
The potential application of Bayesian Bradley-Terry derived rankings, as an alternative for the FIFA ranking or similar systems, represents a promising area for research. Further comparative studies across different sports or competition structures will be conducted to validate the effectiveness of the Bradley-Terry models. As broadly remarked, the interplay between the type of competition, the adopted ranking, and the chosen methodology represents an hot topic for football modellers and deserves a deep and further understanding, both in international matches and domestic leagues.
@@
All analyses were conducted in the R programming language version $4.2.3$ (R Core Team,, 2023). The code to reproduce this manuscript are openly available at https://github.com/RoMaD-96/Bayesian_BTD. The data are available on Kaggle at https://www.kaggle.com/datasets/martj42/international-football-results-from-1872-to-2017.

### v3

@@ 4 Applications
The selection of the training and test sets is crucial and is likely to influence the predictions. We address this by employing an iterative training approach. Our goal-based statistical models and result-based machine learning algorithms are trained on a continuously updated dataset encompassing international matches from $2018$ to $2023$. The matches vary from the FIFA World Cups through the UEFA Euro Championships to normal friendly matches. The data excludes Olympic Games and matches in which at least one of the teams was the national B-team or a U-23 lineup. This approach allows for the incorporation of recent results, potentially improving predictive performance for two distinct scenarios: the group stage and the knockout stage of these two tournaments.
We evaluate the predictive performance of three dynamic goal-based Poisson models, implemented using the footBayes package (Egidi and Palaskas.,, 2022), and three result-based machine learning techniques provided by the caret package (Kuhn,, 2022), both implemented in R, along the lines described in Section 3. Furthermore, we incorporate additional historical information from both the FIFA ranking – through the FIFA ranking points – and the Bayesian BTD derived ranking – through its relative log-strengths – using the bpcs R package (Issa Mattos and Martins Silva Ramos,, 2020).
To ensure a more comparable analysis, both FIFA ranking points and BTD relative log-strengths are normalized using the scaled median absolute deviation (MAD) normalization
@@
where $\mathbb{M}(\cdot)$ is the median. Furthermore, to assess the predictive performance of the models described in Section 3, we employed the Brier score (Brier,, 1950) as recommended by Spiegelhalter and Ng, (2009). It is essentially a mean squared error for forecasts where a lower score indicates greater model predictive accuracy. A common formulation is
@@
The World Cup presents an interesting case study due to the diverse range of National teams with heterogeneous strengths. We specifically investigate the performance of teams in a structured environment, such as the group stage, and contrast it with the dynamic and high-pressure setting of the knockout stage, where single-elimination games can significantly impact teams’ behaviour. This subsection describes how we employ both goal-based statistical models and machine learning algorithms to forecast match outcomes throughout the tournament, assessing their predictive performance by adding the historical information from the ranking systems outlined previously. Notably, we consider the FIFA ranking published just before the World Cup took place, available at https://inside.fifa.com/fifa-world-ranking/men?dateId=id13869.
Figure 1 presents scatterplot between the normalized Bayesian BTD relative log-strengths and the normalized FIFA points. The high value of the Pearson correlation coefficient ($\rho_{p}=0.90$) suggests a positive linear relationship between these two variables. This is further corroborated by the Spearman correlation coefficient ($\rho_{s}=0.88$). Additionally, the Kendall coefficient confirms a substantial positive association ($\tau=0.69$), albeit slightly weaker than the Pearson and Spearman correlations.
Figure 1: 2022 FIFA World Cup. Scatterplot comparing the FIFA ranking points and the BTD relative log-strengths of teams in the World Cup.
@@
The top panel of Figure 2 presents scatterplots comparing the relative strengths of competing teams in both the group and knockout stages under the two ranking systems. Points above the dashed line represent matches where the “away” team had higher relative strength than the “home” team, while points below the line indicate the opposite - we note that the terms ‘home’ and ‘away’ do not mean anything relevant in an international competition, where there are just one or two hosting teams. As expected, both the ranking systems reveal greater variability in team relative strengths during the group stage compared to the knockout stage. This is because teams in the group stage are typically more heterogeneous in terms of strength. As the tournament advances to the knockout stage, the remaining teams become increasingly similar in ability, leading to less variation in relative strength. This pattern is even more evident in the bottom panel of Figure 2, which displays boxplots of relative strength differences between competing teams for each ranking system across the two stages. The boxplots further reveal that the Bayesian BTD ranking system exhibits more variability than the FIFA ranking during the group stage. In contrast, during the knockout stage, relative strength differences variability under both ranking systems are more concentrated, indicating a closer similarity in team abilities.
@@
Figure 2: 2022 FIFA World Cup. The top panel displays scatterplots comparing the normalized relative strengths of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of the normalized relative strength differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two World Cup stages.
@@
The main appeal of these models lies in their ability to predict the outcome of football matches. Table 1 illustrates the predictive accuracy of dynamic Poisson models and machine learning algorithms, measured by the Brier Score, across both group and knockout stages. The goal-based statistical models perform slightly better than the result-based methods during the group stage. In contrast, the machine learning approaches show better predictive accuracy in the knockout stage, where results may be less predictable. A similar trend is seen with ranking systems. While the FIFA ranking system shows marginally better predictive accuracy in the group stage, the Bayesian BTD ranking demonstrates overall better performance during the knockout stage. This suggests that the Bayesian BTD relative log-strengths are particularly apt at predicting outcomes when teams have comparable abilities, which is often the case in the latter stages of the competition.
Model Group Stage Knockout Stage
FIFA BTD FIFA BTD
Diag. Infl. 0.620 0.629 0.530 0.510
Biv. Pois. 0.617 0.618 0.546 0.535
Double Pois. 0.622 0.623 0.543 0.527
MARS 0.640 0.660 0.486 0.503
ANN 0.627 0.660 0.465 0.471
Random Forest 0.713 0.745 0.493 0.461
@@
Table 1: 2022 FIFA World Cup. Brier score for the FIFA ranking and the Bayesian BTD derived ranking across the two World Cup stages.
@@
In this section, we fit the considered statistical models and machine learning algorithms to the data from the most recent CAF Africa Cup of Nations (AFCON) tournament held in Ivory Coast. Notably, the AFCON competition differs from the FIFA World Cup in that the participating teams tend to be more similar in terms of overall strength even during the group stage, making it a compelling case for analyzing the effectiveness of the Bayesian BTD ranking system. In order to conduct this analysis, we use as training set the data from matches played throughout 2018 to the end of 2023. Furthermore, the Bayesian BTD model was executed within this same period to generate team relative log-strengths. In addition, the FIFA ranking employed corresponds to those published on December 21st, available at https://inside.fifa.com/fifa-world-ranking/men?dateId=id14233.
Figure 3 presents a scatterplot for the 2023 CAF Africa Cup of Nations, showing the relationship between normalized Bayesian BTD relative log-strengths and normalized FIFA points. The results are consistent with those from the 2022 World Cup, showing a Pearson correlation coefficient of $\rho_{p}=0.91$, a Spearman correlation coefficient of $\rho_{s}=0.89$, and a Kendall coefficient of $\tau=0.74$, all indicating strong positive associations.
Figure 3: 2023 CAF Africa Cup of Nations. Scatterplot comparing FIFA ranking points and BTD relative log-strengths of teams in the Africa Cup of Nations.
@@
The Bayesian BTD relative log-strengths exhibit less variability compared to the FIFA ranking points, in both the group and knockout stages, as illustrated in the top panel of Figure 4. Furthermore, the boxplots displayed in the bottom panel of Figure 4 show a significant reduction in variance for both ranking systems as we move from the group stage to the knockout stage. However, contrary to what was observed in the World Cup, the Bayesian BTD relative log-strengths exhibit lower variance compared to the FIFA ranking points in the group stage. This may indicate that the Bayesian BTD relative log-strengths more accurately reflect the inherent similarity in team strengths within this tournament stage.
@@
Figure 4: 2023 CAF Africa Cup of Nations. The top panel displays scatterplots comparing the normalized relative strengths of teams in both the group stage (orange dots) and knockout stage (green dots) under the two ranking systems. The dashed grey line represents the bisector. The bottom panel presents boxplots of the normalized relative strength differences for the FIFA ranking (blue) and the Bayesian BTD derived ranking (red) across the two Africa Cup of Nations stages.
@@
As reported in Table 2, all the machine learning algorithms present similar prediction performance in the group stage of the AFCON. However, random forests show the weakest performance in the knockout stage. Furthermore, the other machine learning algorithms and statistical models exhibited similar levels of predictive accuracy in both stages, with statistical models performing slightly better in the group stage. In particular, the inclusion of this alternative ranking system generally improved the predictive performance of most of the statistical models evaluated in the knockout stage.
Model Group Stage Knockout Stage
FIFA BTD FIFA BTD
Diag. Infl. 0.679 0.682 0.681 0.677
Biv. Pois. 0.673 0.682 0.658 0.660
Double Pois. 0.670 0.677 0.670 0.656
MARS 0.679 0.690 0.645 0.650
ANN 0.703 0.702 0.666 0.661
Random Forest 0.736 0.687 0.834 0.884
@@
Table 2: 2023 CAF Africa Cup of Nations. Brier score for the FIFA ranking and the Bayesian BTD derived ranking across the two Africa Cup of Nations stages.
@@ 5 Discussion
This paper investigates the potential improvement in the predictive performance of statistical goal-based methods and machine learning result-based algorithms when a ranking system is incorporated as an additional predictor through its ranking points (relative strengths). We analyze data from the recent 2022 FIFA World Cup in Qatar and the 2023 CAF Africa Cup of Nations in Ivory Coast. Specifically, we explore the effectiveness of a Bayesian Bradley-Terry-Davidson derived ranking system in enhancing prediction accuracy compared to the well-established FIFA ranking system. We compare the performance of these two ranking systems across different tournament stages to identify their potential in predicting match outcomes.
@@
While both the FIFA ranking points and the Bayesian BTD relative log-strengths provide valuable information for predicting outcomes, their effectiveness depends on the stage of the tournament. The FIFA ranking tends to be more accurate during the group stages of the 2022 World Cup, which features more heterogeneous team strengths. Conversely, the Bayesian BTD derived ranking is particularly effective in the group stage of the 2023 AFCON, and it also presents slightly better predictive performances in the knockout stages of both the World Cup and the AFCON. These stages typically exhibit smaller differences in team strengths. Consequently, this result suggests that the Bayesian BTD model effectively captures shifts in team strengths, making it especially valuable in tournaments where competing teams are similar, such as the AFCON, or in stages where the teams exhibit comparable strengths.
@@
It is important to note that while the Bayesian BTD derived ranking is a valuable alternative, it is more computationally intensive compared to the standard FIFA ranking. Specifically, this approach involves a two-step procedure. First, the Bayesian BTD model needs to be computed to derive the teams’ relative log-strengths, and only then their difference can be used as additional predictor in the models being considered.
@@
However, the potential to enhance the accuracy and predictive performances of these models remains significant. Further development could involve refining the Bayesian BTD model to include additional variables that impact match outcomes. These could include the overall market value of the players involved in a specific team, the number of Champions League players, an indicator of the hosting country, or the teams in its neighborhood. Even economic variables such as the GDP per capita or the national population size may be interesting. Furthermore, the implementation of a dynamic methodology, that enables the continuous adjustment for fluctuations in team strength over the course of a season or tournament, could potentially result in enhanced prediction accuracy.
@@
The potential application of Bayesian Bradley-Terry derived rankings, as an alternative for the FIFA ranking or similar systems, represents a promising area for research. Further comparative studies across different sports or competition structures will be conducted to validate the effectiveness of the Bradley-Terry models. As broadly remarked, the interplay between the type of competition, the adopted ranking, and the chosen methodology represents a hot topic for football modellers and deserves a deep and further understanding, both in international matches and domestic leagues.
@@
All analyses were conducted in the R programming language version $4.2.3$ (R Core Team,, 2023). The code to reproduce this manuscript is openly available at https://github.com/RoMaD-96/Bayesian_BTD. The data are available on Kaggle at https://www.kaggle.com/datasets/martj42/international-football-results-from-1872-to-2017.